3. (13 points) Sales, in dollars, for one week were collected for 18 stores in a food store chain. The data is given in the tWill give rating ASAP, please give explanation on how to find the answers with a TI-80 graphing calculator
Answer #1

(a)

The five number summary is,

\small MinValue = 58,762

\small Q1 = \frac{x_4+x_5}{2} = 73,927

\small Median=\frac{x_9+x_10}{2} = 77,605

\small Q3=\frac{x_{14 }+ x_{15}}{2} = 125,983

\small MaxValue=222,585

(b) The box plot is,

| 2300001 | 2200001 | 210000 | 200000 | 190000- | 180000- | 170000- | 160000- | 150000- | 140000- | 130000- | 120000- | 11000

We know that any data point that lies outside the range of [Q1 - 1.5IQR , Q3 + 1.5IQR] is an outlier.

The IQR is 52056

Therefore, the outliers are { 204135 , 222585 }
(c)
The standard deviation is,

\small s=\sqrt\frac{\sum (x-\mu)^2}{n-1}

where, \small \mu = 104,992

Therefore, the standard deviations is,

\small s = 45,327.06

(d)

The distribution is right skewed. As seen from the boxplot, the values stretch far to the right from median then they stretch towards left. Therefore, it can be said that the distribution is skewed to right.